OpenStudy (anonymous):

HELP!! Medal to whoever answers it!!!! What is the equation of the line that passes through the point (3, 2) and has a slope of 4? y = 4x - 1 y = 4x - 5 y = 4x - 10 y = 4x + 1

OpenStudy (anonymous):

so m = 4, x = 3 and y = 2 You need to find the value of y-intercept i.e. "c"" Using above values in y = mx+c, 2 = 4x3 + c on solving, c = -10 Equation = y = mx+c y = 4x-10 = ans.

OpenStudy (nurali):

slope = 4, point (3, 2) (y - 2) = 4(x - 3) y - 2 = 4x - 12 y = 4x -10

OpenStudy (anonymous):

Thank you!!!!! @Nurali @luckythebest